# Consistent Regularization of Signature-Changing BTZ Black Holes

### v1

@@ Abstract
Spacetime singularities pose a fundamental challenge to classical and quantum gravity. We develop a mathematically consistent framework for signature-changing black holes by revisiting the $(2+1)$-dimensional BTZ metric, which undergoes a Lorentzian-to-Euclidean transition at the horizon. We identify and resolve a critical inconsistency in previous regularization schemes concerning second-order distributional terms $\varepsilon^{\prime\prime}(r)$, introducing a modified Hadamard regularization that is rigorously defined within distribution theory. This yields a vacuum solution without surface layers or impulsive waves. Geodesic analysis shows that infalling observers require infinite proper time to reach the horizon, and curvature remains finite everywhere. Furthermore, we establish the physical viability of the solution by proving its linear stability, demonstrating consistent quantum scalar field propagation across the horizon, and reinterpreting the $r=0$ singularity as a removable topological boundary. Our work places signature change and the associated atemporality on a solid foundation as a robust mechanism for singularity resolution.
Signature-change; BTZ black hole; Singularity resolution; Distributional geometry; Hadamard regularization; Quantum gravity
@@ I Introduction
Spacetime singularities, where the predictive power of general relativity breaks down, represent one of the most profound challenges in theoretical physics. While cosmic censorship conjectures often relegate these pathologies behind event horizons Penrose (1969), their existence fundamentally undermines our understanding of spacetime and demands resolution within a complete theory of quantum gravity. Among the various approaches—from string theory Polchinski (1998a, b) and loop quantum gravity Rovelli (2004); Ashtekar and Lewandowski (2004) to non-commutative geometry Szabo (2003)—the paradigm of signature-change offers a particularly radical and elegant mechanism. By allowing the metric signature to transition dynamically from Lorentzian (,+,+)(-,+,+) to Euclidean $(+,+,+)$, this approach can eliminate the causal pathologies that lead to singularities, as exemplified by the Hartle-Hawking no-boundary proposal in cosmology Hartle and Hawking (1983a); Gibbons and Hartle (1990); Ellis and Jaklitsch (1992).
@@
Recently, Capozziello et al. Capozziello et al. (2024) provocatively extended this idea to black hole physics, introducing a signature-changing Schwarzschild metric and proposing atemporality—the cessation of timelike evolution in the interior—as a dynamical mechanism for singularity resolution. While conceptually compelling, their mathematical framework contains a critical flaw: the regularization of distributional curvature relies on manipulations of terms involving $\varepsilon^{\prime\prime}(r)$ that violate the rigorous principles of distribution theory. This inconsistency casts doubt on the validity of their vacuum solution, as it could artificially remove or introduce distributional stress-energy at the change surface.
In this work, we resolve this foundational issue and place the entire paradigm of signature-changing black holes on a solid mathematical footing. We achieve this by developing a consistent geometric framework within the analytically tractable yet physically rich context of the $(2+1)$-dimensional Bañados-Teitelboim-Zanelli (BTZ) black hole Bañados et al. (1992). The BTZ setting provides the ideal theoretical laboratory—complex enough to capture essential black hole features like horizons and curvature, yet simple enough to permit exact computation and unambiguous interpretation.
Our primary contribution is a modified Hadamard regularization scheme that rigorously handles the distributional geometry of signature change, in full compliance with distribution theory. This scheme definitively eliminates the spurious surface layers and impulsive waves that plagued previous approaches, yielding a genuine vacuum solution. Beyond this technical achievement, we provide a comprehensive physical characterization of the resulting geometry, demonstrating through multiple independent lines of evidence that it represents a physically viable and robust resolution of the black hole singularity problem.
The key results and structure of our paper are as follows:
  • In Section II, we construct the signature-changing BTZ metric in both standard and Painlevé-Gullstrand coordinates, and develop our mathematically consistent modified Hadamard regularization scheme. We prove that this yields a distributionally well-defined vacuum geometry ($R_{\mu\nu}=0$, $G_{\mu\nu}=0$) with no delta-sourced stress-energy tensors at the change surface $\Sigma$.
  • Section III provides a comprehensive physical analysis, including the complete regularized curvature tensors showing bounded curvature invariants, the absence of both surface layers and impulsive gravitational waves, and the detailed mechanism of singularity avoidance through atemporality where radially infalling observers require infinite proper time to reach the horizon (see Figs. 3 and 5).
  • Crucially, we go beyond classical consistency to establish physical robustness. In Section IV, we prove the linear stability of the geometry against gravitational perturbations through both analytical and numerical methods (Fig. 7), and demonstrate that quantum scalar field propagation remains well-defined, unitary, and singularity-free across the horizon.
  • We establish the comprehensive physical consistency of the solution by demonstrating that the $r=0$ singularity is resolved into a removable topological boundary, providing a coherent interpretation of the Euclidean interior as a frozen quantum gravitational phase (illustrated in the Penrose diagram, Fig. 6), and proving that standard black hole thermodynamics remains unchanged (Section V).
Our work does more than fix a mathematical inconsistency; it establishes signature change as a serious, well-defined mechanism for singularity resolution in black hole spacetimes. By providing a consistent classical geometry that is regular, stable, and compatible with quantum field propagation, we challenge the necessity of quantum gravity for resolving fundamental spacetime pathologies and open new avenues for exploring the interface between classical and quantum gravity.
@@ II The Signature-Changing BTZ Geometry and its Consistent Regularization
We develop the mathematical foundation for signature-changing black holes by constructing the Lorentzian-Euclidean BTZ metric and establishing a rigorous regularization scheme for its distributional curvature. This provides the consistent geometric framework necessary for analyzing singularity resolution.
@@ IV.2 Quantum Scalar Field Propagation
While the logarithmic term appears mathematically, its physical significance must be assessed through conserved quantities. The conserved current for the scalar field is:
$j^{\mu}=-i(\phi^{*}\nabla^{\mu}\phi-\phi\nabla^{\mu}\phi^{*}).$ (54)
The conservation law $\nabla_{\mu}j^{\mu}=0$ ensures probability preservation. Crucially, the temporal component in our coordinates is:
$j^{\mathscr{T}}=-i\varepsilon(\phi^{*}\partial_{\mathscr{T}}\phi-\phi\partial_{\mathscr{T}}\phi^{*})-i\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}(\phi^{*}\partial_{r}\phi-\phi\partial_{r}\phi^{*}).$ (55)
The factor of $\varepsilon$ in the first term of $j^{\mathscr{T}}$ cancels the singular $\omega^{2}/\varepsilon$ behavior from the equation of motion. The regularized current remains finite and well-defined across the change surface $\Sigma$.
@@
All modes exhibit $\text{Im}(\omega_{n})<0$, confirming exponential decay rather than growth.
@@
The signature-changing BTZ geometry maintains well-defined dynamics for both classical gravitational perturbations and quantum scalar fields. All quasi-normal modes exhibit exponential decay ($\text{Im}(\omega)<0$), confirming linear stability. The current remains conserved, unitarity is preserved, and no information is lost at the signature-changing interface, demonstrating that quantum field theory remains well-defined in our geometry.
@@ IV.3 Physical Interpretation and Comparison
  • Quantum Consistency: Scalar field propagation remains unitary and well-defined Unruh (1976).
@@ V.4 Summary of Physical Consistency
These consistency checks collectively affirm that the paradigm of signature change, when implemented with mathematical rigor, provides a robust and physically plausible mechanism for classical singularity resolution.
@@ VI Discussion and Conclusions
This work has established a mathematically rigorous foundation for signature-changing black holes by resolving critical inconsistencies in previous regularization schemes and demonstrating the viability of atemporality as a mechanism for singularity resolution. Our comprehensive analysis of the $(2+1)$-dimensional BTZ geometry reveals that signature transition at the event horizon provides a natural mechanism for eliminating spacetime singularities without introducing unphysical distributional sources.
@@ VI.2.2 For Quantum Gravity
Our results suggest that some aspects of singularity resolution may be encoded in classical geometry rather than requiring full quantum gravity Rovelli (2004); Ashtekar and Lewandowski (2004). The bounded Kretschmann invariant indicates that curvature need not diverge at the Planck scale, potentially simplifying the interface between classical and quantum gravitational descriptions.
@@ VI.2.3 For Black Hole Physics
VI.3 Future Research Directions
VI.3.1 Immediate Extensions
  • (3+1)-Dimensional Generalization: Application of our regularization scheme to Schwarzschild and Kerr geometries
  • Quantum Field Theory: Investigation of quantum field propagation and vacuum polarization in signature-changing backgrounds
  • Thermodynamic Consistency: Detailed analysis of the third law and phase transitions in signature-changing scenarios
VI.3.2 Conceptual Explorations
  • Connection to Other Approaches: Comparative study with regular black hole models, loop quantum cosmology bounce mechanisms, and string theory fuzzballs
  • Information Paradox: Implications of singularity removal for information preservation and firewall controversies
  • Causal Structure: Deeper investigation of the chronology protection mechanism provided by the change surface
VI.3.3 Observational Signatures
While challenging, potential empirical tests include:
  • Modifications to gravitational wave ringdown signals from the different interior structure
  • Alterations in black hole shadow properties due to matter accumulation near the horizon
  • Effects on accretion disk dynamics and quasi-periodic oscillations
@@ VI.4 Concluding Perspective
The signature-changing paradigm, when placed on firm mathematical footing through consistent regularization, offers a compelling alternative to quantum-gravitational singularity resolution. Our work demonstrates that classical general relativity possesses previously unrecognized capacity for self-regularization through fundamental geometric transformations.

### v3

@@ Abstract
Spacetime singularities represent a fundamental challenge in gravitational physics. We present a mathematically consistent framework for signature-changing black holes based on the $(2+1)$-dimensional BTZ metric, where the signature transitions from Lorentzian (,+,+)(-,+,+) to Euclidean $(+,+,+)$ at the horizon. We identify and rectify a critical inconsistency in previous regularization schemes concerning second-order distributional terms $\varepsilon^{\prime\prime}(r)$, introducing a modified Hadamard regularization that respects distribution theory. This produces a vacuum solution free of surface layers and impulsive gravitational waves. Geodesic analysis reveals that radially infalling observers require infinite proper time to reach the horizon, effectively preventing access to the would-be singularity while maintaining finite curvature invariants throughout the spacetime. We further establish the physical robustness of the geometry by demonstrating linear stability against gravitational perturbations, showing that quantum scalar field propagation remains unitary and well-defined across the signature change, and reinterpreting the $r=0$ region as a topological boundary rather than a curvature singularity. Our work establishes atemporality via signature change as a mathematically rigorous mechanism for singularity avoidance in black hole spacetimes.
Signature-change; BTZ black hole; singularity regularization; Distributional geometry; Hadamard regularization; Quantum gravity
@@ I Introduction
Spacetime singularities, where classical predictability breaks down, represent a central challenge in gravitational physics. While cosmic censorship conjectures typically hide these pathologies behind event horizons Penrose (1969), their existence demands resolution within a consistent framework for quantum gravity. Among various approaches—from string theory Polchinski (1998a, b) and loop quantum gravity Rovelli (2004); Ashtekar and Lewandowski (2004) to noncommutative geometry Szabo (2003)—the paradigm of signature change offers a geometric alternative: by allowing the metric to transition from Lorentzian (,+,+)(-,+,+) to Euclidean $(+,+,+)$, causal pathologies leading to singularities can be avoided, as exemplified in cosmology by the Hartle-Hawking no-boundary proposal Hartle and Hawking (1983a); Gibbons and Hartle (1990); Ellis and Jaklitsch (1992).
@@
Recently, this idea was extended to black hole physics by Capozziello et al. Capozziello et al. (2024), who introduced a signature-changing Schwarzschild metric and proposed atemporality as a mechanism for singularity avoidance. However, their regularization scheme contains a critical mathematical inconsistency: terms involving $\varepsilon^{\prime\prime}(r)$ are manipulated in ways that violate distribution theory, potentially introducing or removing spurious stress-energy at the change surface. This undermines the claim of a vacuum solution and leaves open essential questions of stability, quantum field propagation, and thermodynamic consistency.
In this work, we resolve these issues and establish a mathematically rigorous foundation for signature-changing black holes. Our primary contributions are:
  1. 1.
    A modified Hadamard regularization scheme that respects distribution theory, properly handles $\varepsilon^{\prime\prime}(r)$ terms, and yields a genuine vacuum solution without surface layers or impulsive gravitational waves.
  2. 2.
    A complete physical analysis of the resulting $(2+1)$-dimensional BTZ geometry Bañados et al. (1992), demonstrating bounded curvature, geodesic completeness (infinite proper time to reach the horizon), linear stability, unitary quantum field propagation, and preservation of standard black hole thermodynamics.
  3. 3.
    A clarification of the avoidance mechanism: infinite proper time to cross the horizon constitutes singularity avoidance through atemporality, where the interior becomes Euclidean and causally disconnected from the exterior.
The BTZ setting provides an ideal theoretical laboratory—sufficiently rich to capture essential black hole features, yet tractable enough for exact computation and unambiguous interpretation. We show that signature change, when implemented with mathematical consistency, provides a robust classical mechanism for singularity avoidance, challenging the notion that quantum gravity is strictly necessary for this task.
The paper is structured as follows: Section II presents the signature-changing BTZ metric and our regularization scheme. Section III analyzes curvature and geodesics, demonstrating the atemporality mechanism. Section IV establishes linear stability and unitary quantum field propagation. Section V addresses topological aspects, causal structure, and thermodynamics. We conclude in Section VI with implications and future directions.
@@ II The Signature-Changing BTZ Geometry and its Consistent Regularization
We develop the mathematical foundation for signature-changing black holes by constructing the Lorentzian-Euclidean BTZ metric and establishing a rigorous regularization scheme for its distributional curvature. This provides the consistent geometric framework necessary for analyzing singularity regularization.
@@ IV.2.1 Singularity Structure via Frobenius Method
IV.2.2 Unitarity and Probability Conservation
Despite the logarithmic term in (53), unitarity—the conservation of probability—is preserved. The probability current for a complex scalar field is:
$j^{\mu}=-i\left(\phi^{*}\nabla^{\mu}\phi-\phi\nabla^{\mu}\phi^{*}\right).$ (54)
In our coordinates, the temporal component is:
$j^{\mathscr{T}}=-i\varepsilon\left(\phi^{*}\partial_{\mathscr{T}}\phi-\phi\partial_{\mathscr{T}}\phi^{*}\right)-i\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}\left(\phi^{*}\partial_{r}\phi-\phi\partial_{r}\phi^{*}\right).$ (55)
The crucial observation is that the factor $\varepsilon$ in the first term cancels the singular $\omega^{2}/\varepsilon$ behavior from the equation of motion. To verify finiteness explicitly, substitute the near-horizon expansion $\phi\sim(A+B\ln x)e^{-i\omega\mathscr{T}}$ into (55):
$j^{\mathscr{T}}$ $=-i\varepsilon_{\rho}(x)\left[(A^{*}+B^{*}\ln x)(-i\omega)(A+B\ln x)-(A+B\ln x)(i\omega)(A^{*}+B^{*}\ln x)\right]$
$\quad-i\sqrt{\varepsilon_{\rho}(x)}\sqrt{M-\frac{r_{h}^{2}}{\ell^{2}}}\left[(A^{*}+B^{*}\ln x)\left(\frac{B}{x}+\mathcal{O}(1)\right)-(A+B\ln x)\left(\frac{B^{*}}{x}+\mathcal{O}(1)\right)\right]$
$=2\omega\varepsilon_{\rho}(x)\left[|A|^{2}+2\text{Re}(A^{*}B)\ln x+|B|^{2}(\ln x)^{2}\right]$
$\quad-i\sqrt{\varepsilon_{\rho}(x)}\sqrt{M-\frac{r_{h}^{2}}{\ell^{2}}}\left[\frac{2i\,\text{Im}(A^{*}B)}{x}+\mathcal{O}(\ln x)\right].$ (56)
For $\kappa\geq 2$, $\varepsilon_{\rho}(x)\sim x^{1/(2\kappa+1)}$ vanishes faster than any negative power of $x$, ensuring both terms remain finite. Specifically:
  • First term: $\varepsilon_{\rho}(x)\times\text{polynomial in }\ln x\to 0$ as $x\to 0$
  • Second term: $\sqrt{\varepsilon_{\rho}(x)}/x\sim x^{1/(4\kappa+2)-1}=x^{-(4\kappa+1)/(4\kappa+2)}$ appears singular, but the numerator contains $\text{Im}(A^{*}B)$. By imposing the boundary condition $\text{Im}(A^{*}B)=0$ (which corresponds to choosing a real phase relationship between $A$ and $B$), this term vanishes identically.
The conservation law $\nabla_{\mu}j^{\mu}=0$ can be verified directly from the field equation. Integrating over a spacetime region and applying Gauss’s theorem shows that the total probability $P(\mathscr{T})=\int_{\Sigma_{\mathscr{T}}}\sqrt{h}\,n_{\mu}j^{\mu}\,d^{2}x$ is constant. Thus, despite the logarithmic term in the radial solution, probability is conserved and unitarity is maintained.
IV.2.3 Quasi-normal Mode Spectrum
@@
All modes exhibit $\text{Im}(\omega_{n})<0$, confirming exponential decay rather than growth, consistent with stability.
@@
Quantum scalar field propagation in the signature-changing BTZ geometry is unitary and well-defined. The probability current remains finite and conserved across the change surface $\Sigma$, and all quasi-normal modes exhibit exponential decay ($\text{Im}(\omega)<0$), confirming stability.
@@ IV.3 Physical Interpretation and Comparison
  • Quantum Consistency: Scalar field propagation remains unitary and well-defined across the signature change, as demonstrated by probability current conservation (Section IV.2).
@@ V.4 Summary of Physical Consistency
These consistency checks collectively affirm that the paradigm of signature change, when implemented with mathematical rigor, provides a robust and physically plausible mechanism for classical singularity regularization.
@@ VI Discussion and Conclusions
This work has established a mathematically rigorous foundation for signature-changing black holes by resolving critical inconsistencies in previous regularization schemes and demonstrating the viability of atemporality as a mechanism for singularity avoidance. Our comprehensive analysis of the $(2+1)$-dimensional BTZ geometry reveals that signature transition at the event horizon provides a natural mechanism for eliminating spacetime singularities without introducing unphysical distributional sources.
@@ VI.2.2 For Quantum Gravity
Our results suggest that some aspects of singularity regularization may be encoded in classical geometry rather than requiring full quantum gravity Rovelli (2004); Ashtekar and Lewandowski (2004). The bounded Kretschmann invariant indicates that curvature need not diverge at the Planck scale, potentially simplifying the interface between classical and quantum gravitational descriptions.
@@ VI.2.3 For Black Hole Physics
VI.3 Future Research Directions and Open Questions
Our work establishes a consistent framework for signature-changing black holes in $(2+1)$ dimensions, but several important extensions and open questions remain for future investigation.
VI.3.1 Mathematical and Dimensional Extensions
  • $(3+1)$-Dimensional Generalization: Applying our regularization scheme to Schwarzschild and Kerr geometries presents significant technical challenges due to more complex curvature structure and the presence of genuine curvature singularities. The extension would test whether signature change can similarly regularize four-dimensional singularities.
  • General Signature-Change Formalism: Developing a covariant formulation of signature change that goes beyond the BTZ ansatz, possibly through a smooth scalar field $\Phi$ with $g_{\mu\nu}\partial^{\mu}\Phi\partial^{\nu}\Phi$ changing sign, could provide a more fundamental description.
  • Distributional Geometry Rigor: A more general mathematical treatment of signature-changing metrics within Colombeau algebras or nonlinear distribution theories could further solidify the foundations.
VI.3.2 Physical and Phenomenological Investigations
  • Quantum Field Theory on Signature-Changing Backgrounds: Detailed study of quantum field propagation, vacuum polarization, particle production, and stress-energy tensor regularization in these geometries would clarify their semi-classical consistency.
  • Thermodynamic Refinements: Analysis of the third law, phase transitions between signature-changing and regular black holes, and potential modifications to the entropy-area relation due to the Euclidean interior.
  • Connection to Other Singularity Avoidance Mechanisms: Comparative study with regular black holes, noncommutative geometry, loop quantum gravity bounce mechanisms, and string theory fuzzballs could reveal unifying principles or distinctive features.
VI.3.3 Conceptual and Foundational Questions
  • Information Paradox Implications: If the interior is Euclidean and causally disconnected, how does this affect information retrieval? Does signature change naturally avoid firewalls or promote information conservation?
  • Causal Structure Interpretation: The precise interpretation of the change surface $\Sigma$ as a ”one-way causal membrane” requires deeper analysis within frameworks for non-globally hyperbolic spacetimes and quantum causality.
  • Relation to Quantum Gravity: While our approach demonstrates classical self-regularization, its connection to full quantum gravity—particularly whether the Euclidean interior represents a semiclassical approximation to a genuine quantum gravitational phase—remains open.
VI.3.4 Potential Observational Signatures
While challenging to detect, several potential empirical tests could emerge:
  • Gravitational Wave Modifications: Differences in quasi-normal mode spectra or ringdown behavior due to the modified interior structure, potentially detectable with next-generation gravitational wave observatories.
  • Black Hole Shadow Effects: Matter accumulation near the horizon (due to infinite proper time approach) could subtly alter shadow properties observable with the Event Horizon Telescope and future very-long-baseline interferometry.
  • Accretion Disk Dynamics: Modified interior geometry might influence quasi-periodic oscillations or other accretion phenomena in black hole X-ray binaries.
VI.3.5 Limitations and Caveats
It is important to acknowledge several limitations of our current work:
  • The $(2+1)$-dimensional setting, while mathematically tractable, lacks important physical features of realistic black holes (e.g., dynamical degrees of freedom, realistic matter coupling).
  • The mechanism prevents access to the singularity rather than eliminating it entirely; the $r=0$ region remains a topological boundary, though curvature remains finite.
  • While we demonstrate linear stability, nonlinear stability and long-term evolution remain to be investigated.
These directions highlight that signature change represents a promising but incomplete framework for singularity avoidance. Future work should address these extensions while maintaining the mathematical rigor established here, potentially bridging toward a more comprehensive understanding of spacetime singularities in classical and quantum gravity.
@@ VI.4 Concluding Perspective
The signature-changing paradigm, when placed on firm mathematical footing through consistent regularization, offers a compelling alternative to quantum-gravitational singularity regularization. Our work demonstrates that classical general relativity possesses previously unrecognized capacity for self-regularization through fundamental geometric transformations.